OpenStudy (anonymous):

Solve for x: 8x + 17 = 5(x + 1)

OpenStudy (anonymous):

8X+17 = 5X+5 _5 _5 -------------- 8X+ 12 = 5X -8X -8X --------------- 12= -4X --- ---- -4 -4 -3=X

OpenStudy (anonymous):

x=-4